Burwood Investment Limited has:
1) 8 million shares of common stock outstanding, the price is $32 and beta is 1.15. Market risk premium is 10%, T-bills are yielding 5%.
2) 0.5 million shares of 6% preferred stock outstanding, the price is $67.
3) 100,000 of 9% semi-annual bonds outstanding, 15 years to maturity, and sell for 91% of par. The current yield to maturity on these bonds is 10.18%p.a.
a) If the applicable tax rate is 35% p.a. what is the capital structure of the firm?
b) What is the firm’s WACC?

Please give the calculation process
Answer #1

(1) Number of Common Stock Outstanding = 8 million and Price = $ 32

Value of Common Stock = E = 32 x 8 = $ 256 million

T-Bill Yield = 5 % (Risk-Free Rate), Beta = 1.15 and arket Risk Premium = MRP = 10 %

Using CAPM, the Cost of Equity = Ke = Rf + Beta x MRP = 5 + 1.15 x 10 = 16.5 %

(2) Number of Preferred Shares = 0.5 million, Dividend Yield = 6 % of Par and Current Price = $ 67

Value of Preferred Stock = P = 67 x 0.5 = $ 33.5 million
Preferred Dividend = 0.06 x 100 = $ 6 (assuming a par value of $ 100)

Cost of Preferred Stock = Kp = Dividend / Current Price = 6 / 67 ~ 0.0895 or 8.95 %

(3) Number of Bonds = 100000, Par Value = $ 1000, Selling Price = 91 % of Par = 0.91 x 1000 = $ 910

Value of Bonds = Firm's Debt = D = 100000 x 910 = $ 91000000 = $ 91 million

Cost of Debt = Kd = Yield to maturity of the bonds = 10.18 %

(a) Firms; Capital Structure ;

Common Stock = $ 256 million

Preferred Stock = $ 33,5 million

Debt = $ 91 million

Total = V = $ 380.5 million

(b) Tax Rate = 35 %

WACC = Kd x (1-Tax Rate) x (D/V) + Ke x (E/V) + Kp x (P/V) = 10.18 x (1-0.35) x (91/380.5) + 16.5 x (256/380.5) + 8.95 x (33.5/380.5) = 13.5696 % 13.57 %